# Deblurring 3D Gaussian Splatting

### v1

@@ 4.2 Results and Comparisons
Figure 6: Comparison on applying depth-based pruning. Top: Rendered image from the model with depth-dependent pruning. Bottom: Rendered image from the model with naive pruning as 3D-GS does.
As shown in  table 1 and  fig. 1 our method is on par with the state-of-the-art model in PSNR and achieve state-of-the-art performance evaluated under SSIM on the real defocus dataset. At the same time, it can still enjoy real-time rendering, with noticeable FPS, compared to other deblurring models. Also,  table 2 shows the quantitative result on the synthetic defocus dataset. FPS of 3D-GS is abnormally high because it fails to model the blurry scene, thus only small number of 3D Gaussians are created. fig. 7 and fig. 8 show the qualitative results on real and synthetic defocus dataset, respectively. We can see that ours can produce sharp and fine details, while 3D-GS fail to reconstruct those details.
4.3 Ablation Study
4.3.1 Extra points allocation
@@
In this section, we evaluate the effect of adding extra points to the sparse point cloud. As shown in fig. 5, directly using sparse point cloud without any point densification only represents the objects with a small number of points or totally fails to model the tiny objects. Meanwhile, in case extra points are added to the point cloud, points successfully represent the objects densely. Also, the quantitative result is presented in table 4. It shows assigning valid color features to the additional points is important to deblur the scene and reconstruct the fine details.
@@
4.3.2 Depth-based pruning
We conduct an ablation study on depth-based pruning. To address excessive sparsity of point cloud at the far plane, we prune the points on the far plane less to maintain more numbers of points. table 5 shows there is severe rendering quality degradation if naive pruning, prunes the points with the same threshold from near plane to far plane, is used. In addition, fig. 6 shows a failure to reconstruct objects at the far plane when naive pruning is used, while objects lying on the near-end of the scene are well reconstructed. Meanwhile, ours, with depth-based pruning, can render clean objects on both near and far planes.
Table 3: Transformation of two attributes (scale and rotaion) of 3D Gaussians to model the blur.
train test
scale 0.87 0.85
rotation 0.26 0.25
Methods PSNR $\uparrow$ SSIM $\uparrow$
w/o Extra Points 22.56 0.7016
w/ Random Colors 22.90 0.7089
w/ Extra Points 23.80 0.7430
Methods PSNR $\uparrow$ SSIM $\uparrow$
Naive pruning 23.33 0.7296
Depth-based pruning 23.80 0.7430
5 Limitations & Future Works
NeRF-based deblurring methods [22, 20, 28], which are developed under the assumption of volumetric rendering, are not straightforwardly applicable to rasterization-based 3D-GS [15]. However, they can be compatible to rasterization by optimizing their MLP to deform kernels in the space of the rasterized image instead of letting MLP deform the rays and kernels in the world space. Although it is an interesting direction, it will incur additional costs for interpolating pixels and just implicitly transform the geometry of 3D Gaussians. Therefore, we believe that it will not be an optimal way to model scene blurriness using 3D-GS [15]. Another approach to deblur the defocus blur using 3D-GS [15] can be convolving the rasterized image with conventional grid blur kernels such as gaussian blur kernel [14]. With this approach however, the kernel expressivity is limited compared to the learnable kernels since the Gaussian blur kernel is a simple unimodal Gaussian distribution along each axis and is not learnable. Instead of a fixed Gaussian blur kernel, adopting a learnable grid blur kernel can also be another promising direction to address defocus blur. Nevertheless, as blurriness naturally varies spatially across the scene and each pixel, it is ideal to train blur kernels for every pixel from all training views. Thus, such method can incur high memory requirements, and its usage is limited if the rendering image is high resolution or lots of training images are provided.
6 Conclusion
We proposed Deblurring 3D-GS, the first defocus deblurring algorithm for 3D-GS. We adopted a small MLP that transforms the 3D Gaussians to model the scene blurriness. We also further facilitate deblurring by densifying sparse point clouds from blurry input images through additional point allocation, which uniformly distributes points in the scene and assigns color features using the K-Nearest-Neighbor algorithm. Also, as we applied depth-based pruning instead of naive pruning, which 3D-GS has adopted, we could preserve more points at the far edge of the scene where SfM usually struggles to extract features and fails to generate enough points. Through extensive experiments, we validated that our method can deblur the defocus blur while still enjoying the real-time rendering with FPS $>$ 200. This is because we use the MLP only during the training time, and the MLP is not involved in the inference stage, keeping the inference stage identical to the 3D-GS. Our method achieved state-of-the-art performance or performed comparably with current cutting-edge models evaluated under different metrics.

### v2

@@ B.0.2 Depth-based pruning
Figure 12: Comparison on applying depth-based pruning. Top row: Rendered image from the model with depth-dependent pruning. Bottom row: Rendered image from the model with naive pruning as 3D-GS does.
B.0.3 Ablation study on $M$
We run an ablation study on the hyperparameter $M$, the number of the moments to be averaged. As shown in Tab. 10 our method shows similar performance when $5\leq M\leq 10$. Considering the increasing training time with higher $M$, we set $M$ to 5.
Table 10: Ablation on $M$ under real camera motion blur dataset.
PSNR Training time PSNR Training time
$M=2$ 24.81 $\approx$ 8.0 mins $M=10$ 26.62 $\approx$ 22.4 mins
$M=5$ 26.66 $\approx$ 13.4 mins $M=20$ 25.81 $\approx$ 43.4 mins
B.0.4 Ablation study on extra points allocation
@@
In this section, we evaluate the effect of adding extra points to the sparse point cloud. As shown in Fig. 3, directly using sparse point cloud without any point densification only represents the objects with a small number of points or fails to model the tiny objects. Meanwhile, in case extra points are added to the point cloud, points successfully represent the objects densely. Also, the quantitative result is presented in Tab. 11. It shows assigning valid color features to the additional points is important to deblur the scene and reconstruct the fine details.
Methods PSNR $\uparrow$ SSIM $\uparrow$
w/o Extra Points 23.34 0.7374
w/ Random Colors 23.46 0.7411
w/ Extra Points 23.73 0.7474
B.0.5 Training time comparison
Tab. 12 shows training time and the number of the Gaussians and Tab. 13 describe the running time MLP and rasterization. Although MLP introduces additional costs, our approach is more efficient than prior works that generate kernels and apply convolutions. 3D-GS hardly models the camera motion blur thus fewer 3D Gaussians are involved, dropping the training time. Also, the proposed method uses $M=5$, a relatively smaller number of frames compared to the existing works, leading to faster training time.
Table 12: Comparison on training time (RTX 4090 GPU).
Camera Motion Blur Defocus Blur
Time (min) # Gaussians Time (min) # Gaussians
3D-GS $\approx$ 2.1 63,549 $\approx$ 3.0 308,986
Ours $\approx$ 13.4 139,545 $\approx$ 8.2 317,964
Table 13: Running time of each module.
MLP forwarding 1.33 ms Rasterization 1.00 ms
B.0.6 Images from nearly the same viewpoints
Two training images in Fig. 13 are taken with nearly the same viewpoints but (a) is defocused at the near plane while (b) shows defocus blur at the far plane. Even if they are captured at very close views, the proposed method can deblur only the blurry regions, keeping the clean regions as they are, which highlights its view-dependent functionality.
Figure 13: Deblurred images at nearly the same viewpoints.
@@ B.0.6 Images from nearly the same viewpoints
Figure 16: Qualitative results on synthetic camera motion blur dataset.